# Imitation Learning via Focused Satisficing

### v1

@@ Abstract
Imitation learning often assumes that demonstrations are close to optimal according to some fixed, but unknown, cost function. However, according to satisficing theory, humans often choose acceptable behavior based on their personal (and potentially dynamic) levels of aspiration, rather than achieving (near-) optimality. For example, a lunar lander demonstration that successfully lands without crashing might be acceptable to a novice despite being slow or jerky. Using a margin-based objective to guide deep reinforcement learning, our focused satisficing approach to imitation learning seeks a policy that surpasses the demonstrator’s aspiration levels—defined over trajectories or portions of trajectories—on unseen demonstrations without explicitly learning those aspirations. We show experimentally that this focuses the policy to imitate the highest quality (portions of) demonstrations better than existing imitation learning methods, providing much higher rates of guaranteed acceptability to the demonstrator, and competitive true returns on a range of environments.
@@ 2.3 Subdominance Minimization and Satisficing
with $[x]_{+}\triangleq\max(x,0)$ as the hinge function, and trajectory cost features ${\bf f}:\Xi\rightarrow\mathbb{R}^{K}_{\geq 0}$. Other variants include defining the subdominance using relative cost features, $\text{relsubdom}^{k}_{\alpha_{k}}(\xi,\tilde{\xi})\triangleq\left[\alpha_{k}\left(\frac{f_{k}(\xi)}{f_{k}(\tilde{\xi})}-1\right)+1\right]_{+}$, and/or aggregating over feature dimensions using maximization, $\text{subdom}_{\boldsymbol{\alpha}}(\xi,\tilde{\xi})\triangleq\max_{k}\text{subdom}^{k}_{\alpha_{k}}(\xi,\tilde{\xi})$ Ziebart et al. (2022). Like support vector machines Vapnik and Chapelle (2000), only a subset of support demonstrations, $\tilde{\Xi}_{i}^{\text{SV}_{k}}(\xi_{i})\subseteq\tilde{\Xi_{i}}$, for each task $i$ and feature $k$, actively influence $\boldsymbol{\theta}$:
$\tilde{\xi}_{i,j}\in\tilde{\Xi}_{i}^{\text{SV}_{k}}(\xi_{i})\iff f_{k}(\xi_{i})+\frac{1}{\alpha_{k}}\geq f_{k}(\tilde{\xi}_{i,j}).$ (3)
Unfortunately, optimal control is impractical for many realistic imitation learning problems of interest. Additionally, it makes the learned cost/reward function (Fig. 2) a bottleneck that can prevent the imitation policy from better fitting to (or outperforming) demonstrations.
@@ 2.6 Subdominance Policy Gradient Optimization
The proof for Theorem 4 is provided in our supplementary material. We now present a per state-decomposition of trajectory level subdominance. Subdominance is computed at the final state to determine which features contribute to it, and the contribution of each state-action pair to the total trajectory subdominance is the calculated.
@@
The absolute and relative subdominances for a trajectory $\xi$, from a single task $i$, with respect to a set of demonstrations can be further expanded as:
$\textup{subdom}_{\boldsymbol{\alpha}}(\xi,\tilde{\Xi}_{i})=\!\!\sum_{s_{t}\in\xi,k}\left(\!\frac{\tilde{C}_{i}^{k}}{|\xi|}+\tilde{C}_{i}^{k}\alpha_{k}f_{k}(s_{t})-\frac{\alpha_{k}\tilde{f}_{k,\textup{abs}}^{(i)}}{|\xi||\tilde{\Xi}_{i}|}\right);$
$\textup{relsubdom}_{\boldsymbol{\alpha}}(\xi,\tilde{\Xi}_{i})=\!\!\!\sum_{s_{t}\in\xi,k}\!\left(\frac{\tilde{C}_{i}^{k}(1\!-\!\alpha_{k})}{|\xi|}+\frac{\alpha_{k}f_{k}(s_{t})\tilde{f}_{k,\textup{rel}}^{(i)}}{|\tilde{\Xi}_{i}|}\right),$
where $\tilde{C}_{i}^{k}=\frac{|\tilde{\Xi}_{i}^{\textup{SV}_{k}}|}{|\tilde{\Xi}_{i}|}$, $\tilde{f}_{k,\textup{abs}}^{(i)}=\sum\limits_{\tilde{\xi}\in\tilde{\Xi}_{i}^{\textup{SV}_{k}}}\sum\limits_{s^{\prime}_{t}\in\tilde{\xi}}f_{k}(s^{\prime}_{t})$, and $\tilde{f}_{k,\textup{rel}}^{(i)}=\sum_{\tilde{\xi}\in\tilde{\Xi}_{i}^{\textup{SV}_{k}}}\left(\sum_{s^{\prime}_{t}\in\tilde{\xi}}f_{k}(s^{\prime}_{t})\right)^{-1}$.
This decomposition enables state-of-the-art reinforcement learning algorithms Schulman et al. (2017) that assign credit to actions in a causally consistent manner (i.e., only future returns influence an action’s updates) to be employed. Further flexibility is gained via the choice of policy representation. The derivation is provided in Appendix C.
@@ 2.7 Subdominance Policy Gradient Algorithms
Algorithm 1 outlines our approach for optimization. For each task ($i$), a trajectory is rolled out by sampling from the current learned policy (Line 2). The cost features of the sampled trajectory and the demonstrated trajectory are compared to determine which dimensions the sampled trajectory does not sufficiently outperform the demonstration, and are thus support vectors (Line 4). Here, the $\alpha$ values defining margin slopes (Eq. (3)) can either be optimized numerically (e.g., using stochastic gradient descent) or analytically Memarrast et al. (2023) as described in Appendix A. A policy update is then employed to reduce the subdominance (Line 7).
@@
$\boldsymbol{\theta}\!\leftarrow\!\boldsymbol{\theta}\!+\!\eta\sum_{\mathclap{i,\tilde{\xi}_{i,j}\in\tilde{\Xi}_{i}}}\tilde{r}_{\theta}^{(i,j)}\textup{subdom}_{\boldsymbol{\alpha}}(\tilde{\xi}_{i,j},\tilde{\Xi}_{i})\sum_{\mathclap{(s,a)\in\tilde{\xi}_{i,j}}}\nabla_{\boldsymbol{\theta}}\!\log\pi_{\boldsymbol{\theta}}\!\left(a|s\right),$ (10)
where $\tilde{r}_{\theta}^{(i,j)}=\frac{\pi_{\boldsymbol{\theta}}(\tilde{\xi}_{i,j})}{\tilde{\pi}(\tilde{\xi}_{i,j})}$ is the importance ratio, and $\tilde{\pi}$ is an estimate of the demonstrator’s policy.
@@
1:  Estimate $\tilde{\pi}$ by performing behavior cloning on demonstrations $\tilde{\Xi}$
@@
6:           $\alpha_{k}\!\leftarrow\!\alpha_{k}\exp\bigl{\{}-{\eta}_{t}^{\prime}\tilde{r}_{\theta}^{(i,j)}\sum\limits_{\mathclap{\tilde{\xi}_{i,m}\in\tilde{\Xi}_{i,j}^{\text{SV}_{k}}(\alpha_{k})}}\bigl{(}f_{k}(\tilde{\xi}_{i,j})\!\!-\!f_{k}(\tilde{\xi}_{i,m})\bigr{)}\!+\!\lambda|\tilde{\Xi}|\alpha_{k}\bigr{\}}$
@@ 3.3 Training and Bootstrapping
Using Algorithm 1 and analytically computed $\alpha$ values in step 4, we initialize our Online MinSubFI training with a policy that is pretrained via Offline MinSubFI (Corollary 10); we motivate this choice via an ablation study with different policy initializations in Appendix F. For all of our experiments throughout the paper, we employ a quadratic expansion of the original cost features, as described in Appendix E.2.
@@
Baselines Ours
Environment Demonstrations BC TREX TREXCF AIRL GAIL MinSubFIOFF MinSubFION MinSubFISNIP MinSubFISNIP* MinSubFILCF
@@ 3.7 Performance with Human Demonstrations
4 Conclusions & Future Work
@@ Acknowledgments
Appendix A Optimization of $\boldsymbol{\alpha}$: Numerical and Analytical
@@ Appendix C Per-State Cost Decomposition
$\textup{subdom}^{\Sigma}_{\boldsymbol{\alpha}}(\xi,\tilde{\Xi}_{i})=\sum_{k}\text{subdom}^{k}_{\alpha_{k}}(\xi,\tilde{\Xi}_{i})$
$=\sum_{k}\frac{1}{|\tilde{\Xi}_{i}|}\sum_{\tilde{\xi}\in\tilde{\Xi}_{i}}\left[\alpha_{k}(f_{k}(\xi)-f_{k}(\tilde{\xi}))+1\right]_{+}$
$=\sum_{k}\frac{1}{|\tilde{\Xi}_{i}|}\sum_{\tilde{\xi}\in\tilde{\Xi}_{i}^{\textup{SV}_{k}}}\left(\alpha_{k}(f_{k}(\xi)-f_{k}(\tilde{\xi}))+1\right)$ (15)
$=\sum_{k}\frac{\alpha_{k}}{|\tilde{\Xi}_{i}|}\sum_{{\tilde{\xi}\in\tilde{\Xi}_{i}^{\textup{SV}_{k}}}}\left(\frac{1}{\alpha_{k}}+\sum_{\mathclap{s_{t}\in\xi}}f_{k}(s_{t})-\sum_{\mathclap{s^{\prime}_{t}\in\tilde{\xi}}}f_{k}(s^{\prime}_{t})\right)$
$=\sum_{s_{t}\in\xi,k}\frac{\alpha_{k}}{|\tilde{\Xi}_{i}|}\sum_{\tilde{\xi}\in\tilde{\Xi}_{i}^{\textup{SV}_{k}}}\left(\frac{1}{\alpha_{k}|\xi_{i}|}+f_{k}(s_{t})-\frac{\sum\limits_{\mathclap{s^{\prime}_{t}\in\tilde{\xi}}}f_{k}(s^{\prime}_{t})}{|\xi|}\right)$
$=\sum_{\mathclap{s_{t}\in\xi,k}}\frac{\alpha_{k}|\tilde{\Xi}_{i}^{\textup{SV}_{k}}|}{|\tilde{\Xi}_{i}|}\!\left(\frac{1}{\alpha_{k}|\xi|}\!+\!f_{k}(s_{t})\!-\!\frac{\sum\limits_{\tilde{\xi}\in\tilde{\Xi}_{i,n}^{\textup{SV}_{k}}}\sum\limits_{s^{\prime}_{t}\in\tilde{\xi}}f_{k}(s^{\prime}_{t})}{|\xi_{i,n}||\tilde{\Xi}_{i,n}^{\textup{SV}_{k}}|}\right)$
$=\sum_{s_{t}\in\xi_{i},k}\frac{\tilde{C}_{i}^{k}}{|\xi_{i}|}+\tilde{C}_{i}^{k}\alpha_{k}f_{k}(s_{t})-\frac{\alpha_{k}\tilde{f}_{k,\textup{abs}}^{(i)}}{|\xi_{i}||\tilde{\Xi}_{i}|},$
where $\tilde{C}_{i}^{k}=\frac{|\tilde{\Xi}_{i}^{\textup{SV}_{k}}|}{|\tilde{\Xi}_{i}|}$, and $\tilde{f}_{k,\textup{abs}}^{(i)}=\sum\limits_{\tilde{\xi}\in\tilde{\Xi}_{i}^{\textup{SV}_{k}}}\sum\limits_{s^{\prime}_{t}\in\tilde{\xi}}f_{k}(s^{\prime}_{t})$. ∎
@@
$\textup{relsubdom}^{\Sigma}_{\boldsymbol{\alpha}}(\xi,\tilde{\Xi}_{i})=\sum_{k}\text{relsubdom}^{k}_{\alpha_{k}}(\xi,\tilde{\Xi}_{i})$
$=\sum_{k}\frac{1}{|\tilde{\Xi}_{i}|}\sum_{\tilde{\xi}\in\tilde{\Xi}_{i}}\left[\alpha_{k}\left(\frac{f_{k}(\xi)}{f_{k}(\tilde{\xi})}-1\right)+1\right]_{+}$
$=\sum_{k}\frac{1}{|\tilde{\Xi}_{i}|}\sum_{\tilde{\xi}\in\tilde{\Xi}_{i}^{\textup{SV}_{k}}}\left(\alpha_{k}\left(\frac{f_{k}(\xi)}{f_{k}(\tilde{\xi})}-1\right)+1\right)$ (16)
$=\sum_{k}\frac{1}{|\tilde{\Xi}_{i}|}\sum_{\tilde{\xi}\in\tilde{\Xi}_{i}^{\textup{SV}_{k}}}\left((\beta_{k}-\alpha_{k})+\alpha_{k}\frac{\sum_{s_{t}\in\xi}f_{k}(s_{t})}{\sum_{s^{\prime}_{t}\in\tilde{\xi}}f_{k}(s^{\prime}_{t})}\right)$
$=\sum_{s_{t}\in\xi,k}\!\Biggl{(}\frac{(1-\alpha_{k})|\tilde{\Xi}_{i}^{\textup{SV}_{k}}|}{|\xi_{i}||\tilde{\Xi}_{i}|}\!+\!\frac{\alpha_{k}f_{k}(s_{t})}{|\tilde{\Xi}_{i}|}\!\sum\limits_{\tilde{\xi}\in\tilde{\Xi}_{i}^{\textup{SV}_{k}}}\frac{1}{\sum\limits_{\mathclap{s^{\prime}_{t}\in\tilde{\xi}}}f_{k}(s^{\prime}_{t})}\Biggr{)}$
$=\sum_{s_{t}\in\xi_{i},k}\frac{\tilde{C}_{i}^{k}(1-\alpha_{k})}{|\xi_{i}|}+\frac{\alpha_{k}f_{k}(s_{t})\tilde{f}_{k,\textup{rel}}^{(i)}}{|\tilde{\Xi}_{i}|},$
where $\tilde{C}_{i}^{k}=\frac{|\tilde{\Xi}_{i}^{\textup{SV}_{k}}|}{|\tilde{\Xi}_{i}|}$, and $\tilde{f}_{k,\textup{rel}}^{(i)}=\sum\limits_{\tilde{\xi}\in\tilde{\Xi}_{i}^{\textup{SV}_{k}}}\left(\sum\limits_{s^{\prime}_{t}\in\tilde{\xi}}f_{k}(s^{\prime}_{t})\right)^{-1}$. ∎

### v2

@@ Abstract
Imitation learning often assumes that demonstrations are close to optimal according to some fixed, but unknown, cost function. However, according to satisficing theory, humans often choose acceptable behavior based on their personal (and potentially dynamic) levels of aspiration, rather than achieving (near-) optimality. For example, a lunar lander demonstration that successfully lands without crashing might be acceptable to a novice despite being slow or jerky. Using a margin-based objective to guide deep reinforcement learning, our focused satisficing approach to imitation learning seeks a policy that surpasses the demonstrator’s aspiration levels—defined over trajectories or portions of trajectories—on unseen demonstrations without explicitly learning those aspirations. We show experimentally that this focuses the policy to imitate the highest quality (portions of) demonstrations better than existing imitation learning methods, providing much higher rates of guaranteed acceptability to the demonstrator, and competitive true returns on a range of environments.Accepted for publication at the 34th International Joint Conference on Artificial Intelligence (IJCAI 2025).
@@ 2.3 Subdominance Minimization and Satisficing
with $[x]_{+}\triangleq\max(x,0)$ as the hinge function, and trajectory cost features ${\bf f}:\Xi\rightarrow\mathbb{R}^{K}_{\geq 0}$. Other variants include defining the subdominance using relative cost features, $\text{relsubdom}^{k}_{\alpha_{k}}(\xi,\tilde{\xi})\triangleq\left[\alpha_{k}\left(\frac{f_{k}(\xi)}{f_{k}(\tilde{\xi})}-1\right)+1\right]_{+}$, and/or aggregating over feature dimensions using maximization, $\text{subdom}_{\boldsymbol{\alpha}}(\xi,\tilde{\xi})\triangleq\max_{k}\text{subdom}^{k}_{\alpha_{k}}(\xi,\tilde{\xi})$ Ziebart et al. (2022). Like support vector machines Vapnik and Chapelle (2000), only a subset of support demonstrations, $\tilde{\Xi}_{i}^{\text{SV}_{k}}(\xi)\subseteq\tilde{\Xi_{i}}$, for each task $i$ and feature $k$, actively influence $\boldsymbol{\theta}$:
$\tilde{\xi}\in\tilde{\Xi}_{i}^{\text{SV}_{k}}(\xi)\iff f_{k}(\xi)+\frac{1}{\alpha_{k}}\geq f_{k}(\tilde{\xi}).$ (3)
For notational convenience, when $\xi$ is indexed (e.g., by $(i,j)$ as $\xi_{i,j}$), we denote this resulting support vector set for all demonstrations of task $i$ as $\tilde{\Xi}_{i,j}^{\text{SV}_{k}}$ for feature $k$. Unfortunately, optimal control is impractical for many realistic imitation learning problems of interest. Additionally, it makes the learned cost/reward function (Fig. 2) a bottleneck that can prevent the imitation policy from better fitting to (or outperforming) demonstrations.
@@ 2.6 Subdominance Policy Gradient Optimization
The proof for Theorem 4 is provided in our supplementary material. We now present a state-based decomposition of trajectory level subdominance. Subdominance is computed at the final state to determine which features contribute to it, and the contribution of each state-action pair to the total trajectory subdominance is the calculated.
@@
The absolute and relative subdominances for a trajectory $\xi$, with respect to a set of demonstrations $\Xi$ can be further expanded as:
$\textup{subdom}_{\boldsymbol{\alpha}}(\xi,\tilde{\Xi})=\!\!\sum_{s_{t}\in\xi,k}\left(\!\frac{\tilde{C}_{\xi,\tilde{\Xi}}^{k}}{|\xi|}+\tilde{C}_{\xi,\tilde{\Xi}}^{k}\alpha_{k}f_{k}(s_{t})-\frac{\alpha_{k}\tilde{f}_{k,\xi,\tilde{\Xi}}^{\textup{abs}}}{|\xi||\tilde{\Xi}|}\right);$
$\textup{relsubdom}_{\boldsymbol{\alpha}}(\xi,\tilde{\Xi})=\!\!\!\sum_{s_{t}\in\xi,k}\!\left(\frac{\tilde{C}_{\xi,\tilde{\Xi}}^{k}(1\!-\!\alpha_{k})}{|\xi|}+\frac{\alpha_{k}f_{k}(s_{t})\tilde{f}_{k,\xi,\tilde{\Xi}}^{\textup{rel}}}{|\tilde{\Xi}|}\right),$
where $\tilde{C}_{\xi,\tilde{\Xi}}^{k}=\frac{|\tilde{\Xi}^{\textup{SV}_{k}}(\xi)|}{|\tilde{\Xi}|}$, $\tilde{f}_{k,\xi,\tilde{\Xi}}^{\textup{abs}}=\sum\limits_{\tilde{\xi}\in\tilde{\Xi}^{\textup{SV}_{k}}(\xi)}\sum\limits_{s^{\prime}_{t}\in\tilde{\xi}}f_{k}(s^{\prime}_{t})$, and $\tilde{f}_{k,\xi,\tilde{\Xi}}^{\textup{rel}}=\sum_{\tilde{\xi}\in\tilde{\Xi}^{\textup{SV}_{k}}(\xi)}\left(\sum_{s^{\prime}_{t}\in\tilde{\xi}}f_{k}(s^{\prime}_{t})\right)^{-1}$.
This decomposition enables state-of-the-art reinforcement learning algorithms Schulman et al. (2017) that assign credit to actions in a causally consistent manner (i.e., only future returns influence an action’s updates) to be employed. Further flexibility is gained via the choice of policy representation.
@@ 2.7 Subdominance Policy Gradient Algorithms
Algorithm 1 outlines our approach for optimization. For each task ($i$), a trajectory is rolled out by sampling from the current learned policy (Line 2). The cost features of the sampled trajectory and the demonstrated trajectory are compared to determine which dimensions the sampled trajectory does not sufficiently outperform the demonstration, and are thus support vectors (Line 4). Here, the $\alpha$ values defining margin slopes (Eq. (3)) can either be optimized numerically (e.g., using stochastic gradient descent) or analytically Memarrast et al. (2023). A policy update is then employed to reduce the subdominance (Line 7).
@@
$\boldsymbol{\theta}\!\leftarrow\!\boldsymbol{\theta}\!+\!\eta\sum_{\mathclap{i,\tilde{\xi}_{i,j}\in\tilde{\Xi}_{i}}}\tilde{r}_{\theta,\tilde{\pi}}^{(i,j)}\textup{subdom}_{\boldsymbol{\alpha}}(\tilde{\xi}_{i,j},\tilde{\Xi}_{i})\sum_{\mathclap{(s,a)\in\tilde{\xi}_{i,j}}}\nabla_{\boldsymbol{\theta}}\!\log\pi_{\boldsymbol{\theta}}\!\left(a|s\right),$ (10)
where $\tilde{r}_{\theta,\tilde{\pi}}^{(i,j)}=\frac{\pi_{\boldsymbol{\theta}}(\tilde{\xi}_{i,j})}{\tilde{\pi}(\tilde{\xi}_{i,j})}$ is the importance ratio, and $\tilde{\pi}$ is an estimate of the demonstrator’s policy.
@@
1:  Estimate $\tilde{\pi}_{\text{BC}}$ using behavior cloning on demonstrations $\tilde{\Xi}$
@@
6:           $\alpha_{k}\!\leftarrow\!\alpha_{k}\exp\bigl{\{}-{\eta}_{t}^{\prime}\tilde{r}_{\theta,\tilde{\pi}_{\text{BC}}}^{(i,j)}\sum\limits_{\mathclap{\tilde{\xi}^{\prime}\in\tilde{\Xi}_{i,j}^{\text{SV}_{k}}}}\bigl{(}f_{k}(\tilde{\xi}_{i,j})\!-\!f_{k}(\tilde{\xi}^{\prime})\bigr{)}\!+\!\lambda|\tilde{\Xi}|\alpha_{k}\bigr{\}}$
@@ 3.3 Training and Bootstrapping
Using Algorithm 1 and analytically computed $\alpha$ values in step 4, we initialize our Online MinSubFI training with a policy that is pretrained via Offline MinSubFI (Corollary 10); we motivate this choice via an ablation study with different policy initializations in the extended version of this paper. For all of our experiments throughout the paper, we employ a quadratic expansion of the original cost features as the vectorized outer product of the original cost feature vector, ${\bf f}_{\text{expanded}}=\text{vec}({\bf f}\cdot{\bf f}^{\top})$.
@@
Demon- Baselines Ours
Environment strations BC TREX TREXCF AIRL GAIL MinSubFIOFF MinSubFION MinSubFISNIP MinSubFISNIP* MinSubFILCF
@@ 3.7 Performance with Human Demonstrations
4 Conclusions and Future Work
@@ Acknowledgments
Appendix A Optimization of ${\alpha}$: Numerical and Analytical
@@ Appendix C Per-State Cost Decomposition
$\textup{subdom}^{\Sigma}_{\boldsymbol{\alpha}}(\xi,\tilde{\Xi})=\sum_{k}\text{subdom}^{k}_{\alpha_{k}}(\xi,\tilde{\Xi})$
$=\sum_{k}\frac{1}{|\tilde{\Xi}|}\sum_{\tilde{\xi}\in\tilde{\Xi}}\left[\alpha_{k}(f_{k}(\xi)-f_{k}(\tilde{\xi}))+1\right]_{+}$
$=\sum_{k}\frac{1}{|\tilde{\Xi}|}\sum_{\tilde{\xi}\in\tilde{\Xi}^{\textup{SV}_{k}}(\xi)}\left(\alpha_{k}(f_{k}(\xi)-f_{k}(\tilde{\xi}))+1\right)$ (15)
$=\sum_{k}\frac{\alpha_{k}}{|\tilde{\Xi}|}\sum_{{\tilde{\xi}\in\tilde{\Xi}^{\textup{SV}_{k}}(\xi)}}\left(\frac{1}{\alpha_{k}}+\sum_{\mathclap{s_{t}\in\xi}}f_{k}(s_{t})-\sum_{\mathclap{s^{\prime}_{t}\in\tilde{\xi}}}f_{k}(s^{\prime}_{t})\right)$
$=\sum_{s_{t}\in\xi,k}\frac{\alpha_{k}}{|\tilde{\Xi}|}\sum_{\tilde{\xi}\in\tilde{\Xi}^{\textup{SV}_{k}}(\xi)}\left(\frac{1}{\alpha_{k}|\xi|}+f_{k}(s_{t})-\frac{\sum\limits_{\mathclap{s^{\prime}_{t}\in\tilde{\xi}}}f_{k}(s^{\prime}_{t})}{|\xi|}\right)$
$=\sum_{\mathclap{s_{t}\in\xi,k}}\frac{\alpha_{k}|\tilde{\Xi}^{\textup{SV}_{k}}(\xi)|}{|\tilde{\Xi}|}\!\left(\frac{1}{\alpha_{k}|\xi|}\!+\!f_{k}(s_{t})\!-\!\frac{\sum\limits_{\tilde{\xi}\in\tilde{\Xi}^{\textup{SV}_{k}}(\xi)}\sum\limits_{s^{\prime}_{t}\in\tilde{\xi}}f_{k}(s^{\prime}_{t})}{|\xi||\tilde{\Xi}^{\textup{SV}_{k}}(\xi)|}\right)$
$=\sum_{s_{t}\in\xi,k}\frac{\tilde{C}_{\xi,\tilde{\Xi}}^{k}}{|\xi|}+\tilde{C}_{\xi,\tilde{\Xi}}^{k}\alpha_{k}f_{k}(s_{t})-\frac{\alpha_{k}\tilde{f}_{k,\xi,\tilde{\Xi}}^{\textup{abs}}}{|\xi||\tilde{\Xi}|},$
where $\tilde{C}_{\xi,\tilde{\Xi}}^{k}=\frac{|\tilde{\Xi}^{\textup{SV}_{k}}(\xi)|}{|\tilde{\Xi}|}$, and $\tilde{f}_{k,\xi,\tilde{\Xi}}^{\textup{abs}}=\sum\limits_{\tilde{\xi}\in\tilde{\Xi}^{\textup{SV}_{k}}(\xi)}\sum\limits_{s^{\prime}_{t}\in\tilde{\xi}}f_{k}(s^{\prime}_{t})$.
@@
$\textup{relsubdom}^{\Sigma}_{\boldsymbol{\alpha}}(\xi,\tilde{\Xi})=\sum_{k}\text{relsubdom}^{k}_{\alpha_{k}}(\xi,\tilde{\Xi}_{i})$
$=\sum_{k}\frac{1}{|\tilde{\Xi}|}\sum_{\tilde{\xi}\in\tilde{\Xi}}\left[\alpha_{k}\left(\frac{f_{k}(\xi)}{f_{k}(\tilde{\xi})}-1\right)+1\right]_{+}$
$=\sum_{k}\frac{1}{|\tilde{\Xi}|}\sum_{\tilde{\xi}\in\tilde{\Xi}^{\textup{SV}_{k}}(\xi)}\left(\alpha_{k}\left(\frac{f_{k}(\xi)}{f_{k}(\tilde{\xi})}-1\right)+1\right)$ (16)
$=\sum_{k}\frac{1}{|\tilde{\Xi}|}\sum_{\tilde{\xi}\in\tilde{\Xi}^{\textup{SV}_{k}}(\xi)}\left((\beta_{k}-\alpha_{k})+\alpha_{k}\frac{\sum_{s_{t}\in\xi}f_{k}(s_{t})}{\sum_{s^{\prime}_{t}\in\tilde{\xi}}f_{k}(s^{\prime}_{t})}\right)$
$=\sum_{s_{t}\in\xi,k}\!\Biggl{(}\frac{(1-\alpha_{k})|\tilde{\Xi}^{\textup{SV}_{k}}(\xi)|}{|\xi||\tilde{\Xi}|}\!+\!\frac{\alpha_{k}f_{k}(s_{t})}{|\tilde{\Xi}|}\!\sum\limits_{\tilde{\xi}\in\tilde{\Xi}^{\textup{SV}_{k}}(\xi)}\frac{1}{\sum\limits_{\mathclap{s^{\prime}_{t}\in\tilde{\xi}}}f_{k}(s^{\prime}_{t})}\Biggr{)}$
$=\sum_{s_{t}\in\xi,k}\frac{\tilde{C}_{\xi,\tilde{\Xi}}^{k}(1-\alpha_{k})}{|\xi|}+\frac{\alpha_{k}f_{k}(s_{t})\tilde{f}_{k,\xi,\tilde{\Xi}}^{\textup{rel}}}{|\tilde{\Xi}|},$
where $\tilde{C}_{\xi,\tilde{\Xi}}^{k}=\frac{|\tilde{\Xi}^{\textup{SV}_{k}}|}{|\tilde{\Xi}_{i}|}$, and $\tilde{f}_{k,\xi,\tilde{\Xi}}^{\textup{rel}}=\sum\limits_{\tilde{\xi}\in\tilde{\Xi}^{\textup{SV}_{k}}(\xi)}\left(\sum\limits_{s^{\prime}_{t}\in\tilde{\xi}}f_{k}(s^{\prime}_{t})\right)^{-1}$. ∎